OpenStudy (anonymous):

Situation: A 11 gram sample of a substance that's used to treat thyroid disorders has a k-value of .125. find the substance's half-life, in days. round to the nearest tenth

OpenStudy (anonymous):

.125^t=.5 find t

OpenStudy (anonymous):

how do i find t ?

OpenStudy (anonymous):

log form

OpenStudy (anonymous):

explain? or show me?

OpenStudy (anonymous):

\[\log_ab=c\] can be written as a^c=b

OpenStudy (anonymous):

your formula is .125^t=.5 so \[\log_{.125}.5=t\]

OpenStudy (anonymous):

1.5 ?

OpenStudy (anonymous):

how's you get that?

OpenStudy (anonymous):

is .5 a exponent?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i just got 1.4 now

OpenStudy (anonymous):

log .5/ log .125

OpenStudy (anonymous):

.33 is what i just got

OpenStudy (anonymous):

that's right